### v1

# Code-Optimise: Self-Generated Preference Data for Correctness and Efficiency

Abstract
Code Language Models have been trained to generate accurate solutions, typically with no regard for runtime. On the other hand, previous works that explored execution optimisation have observed corresponding drops in functional correctness. To that end, we introduce Code-Optimise, a framework that incorporates both correctness (passed, failed) and runtime (quick, slow) as learning signals via self-generated preference data. Our framework is both lightweight and robust as it dynamically selects solutions to reduce overfitting while avoiding a reliance on larger models for learning signals. Code-Optimise achieves significant improvements in $pass@k$ while decreasing the competitive baseline runtimes by an additional 6% for in-domain data and up to 3% for out-of-domain data. As a byproduct, the average length of the generated solutions is reduced by up to 48% on MBPP and 23% on HumanEval, resulting in faster and cheaper inference. The generated data and codebase will be open-sourced at www.open-source.link.
Code-Optimise: Self-Generated Preference Data
for Correctness and Efficiency
Leonidas Gee1thanks:  Research internship at Huawei Noah’s Ark Lab, London., Milan Gritta2, Gerasimos Lampouras2, and Ignacio Iacobacci2 1University of Sussex, Brighton, UK 2Huawei Noah’s Ark Lab, London, UK jg717@sussex.ac.uk {milan.gritta, gerasimos.lampouras, ignacio.iacobacci}@huawei.com
Figure 1: Overview of Code-Optimise. (1) Diverse solutions are sampled per problem. (2) A code interpreter annotates the solutions by functional correctness and runtime. (3) CLM is optimised using SFT or DPO.
Model Split Problem Solution
Total Filtered Ratio Total Filtered Ratio CoV
StarCoder-1B Train 384 183 47.66 38400 15472 40.29 0.011
Validation 90 40 44.44 9000 3533 39.26 0.010
StarCoder-3B Train 384 211 54.95 38400 17575 45.77 0.007
Validation 90 45 50.00 9000 3926 43.62 0.014
CodeLlama-7B Train 384 250 65.10 38400 21350 55.60 0.007
Validation 90 55 61.11 9000 4962 55.13 0.008
CodeLlama-13B Train 384 261 67.97 38400 22182 57.77 0.007
Validation 90 56 62.22 9000 5108 56.76 0.007
Table 1: Statistics of our self-generated preference data. 1) A Model generates 100 solutions per problem out of Total problems in each Split. 2) Functional correctness and runtime are annotated. 3) Problems are filtered to retain those with at least 2 passing and 1 failing solution (Filtered). A low coefficient of variation (CoV $\leq$ 0.1) across 5 runs indicates that runtime measurements are stable. Ratio is the percentage of $\frac{\textbf{Filtered}}{\textbf{Total}}$ retained code solutions.
1 Introduction
Code Language Models (CLMs) trained on large code repositories such as The Stack Kocetkov et al. (2022); Lozhkov et al. (2024) gradually increase their understanding of code semantics. CLMs are thus able to generate functionally correct and reasonably efficient solutions to programming problems Austin et al. (2021); Chen et al. (2021), among many other code related skills Li et al. (2023). Shypula et al. (2023) have shown that CLMs can optimise slow-running code to achieve large runtime gains but at a substantial cost to correctness (down by up to $\sim$30%). Subsequent research has focused mostly on improving code correctness. On the data perspective, a common way of improving functional correctness is via distilled supervised fine-tuning Tunstall et al. (2023); Xu et al. (2023); Luo et al. (2023); Wei et al. (2023) on training data generated by large models such as GPT-4 Achiam et al. (2023). However, in many cases, due to legal, financial and/or privacy constraints, it is not feasible to rely on proprietary data. Additionally, we seek to overcome the limitations of supervised fine-tuning (SFT), which only optimises for ‘positive’ examples, with no means of reducing the likelihood of generating undesirable (e.g. incorrect or slow) code. Although such issues may be addressed via Reinforcement Learning (RL) Le et al. (2022); Wang et al. (2022); Gorinski et al. (2023), they often come with added complexity and instability. Therefore, we opt for Direct Preference Optimisation Rafailov et al. (2024) as our preferred fine-tuning method due to its simplicity and widespread adoption. We propose Code-Optimise, a lightweight framework that trains CLMs with our self-generated preference data for correctness (passed/failed) and efficiency (quick/slow), shown in Figure 1. Starting from a small collection of problems and unit tests, Code-Optimise bootstraps the pretrained CLM to generate the required learning signals thereby exposing the model to on-policy automatically annotated data. Code-Optimise provides additional robustness by dynamically selecting solutions during training to reduce overfitting. Our method consists of three steps: 1) Sampling; generate N solutions for each problem description, 2) Annotation; automatically label each solution for correctness and runtime, 3) Optimisation; fine-tune the CLM on the self-generated preference data using several lightweight configurations. The main contributions of Code-Optimise are:
  • We create and publish a novel code preference dataset (and a recipe to extend it) that enables multi-objective optimisation (code correctness and runtime efficiency) of CLMs.
  • We present experimental analysis to support our approach and observe that functional correctness is significantly improved, particularly for smaller CLMs and lower $k$ in $pass@k$. The scores are further enhanced with our Dynamic Solution Selection (DSS).
  • We demonstrate that runtimes are reduced by up to 6% for MBPP and up to 3% for HumanEval over competitive baseline CLMs. Finally, the length of generated solutions is reduced by up to 23% for HumanEval and up to 48% for MBPP, decreasing inference costs.
To the best of our knowledge, our work is the first to show improvements in both correctness and efficiency for the task of code generation.
2 Code-Optimise
We now introduce Code-Optimise, a lightweight method for optimisation of CLMs aimed at improving functional correctness of code as well as reducing its runtime, shown in Figure 1.
2.1 Sampling
We assume access to $D_{seed}=\{x_{i},y_{i},ut_{i}\}_{i=1}^{N}$, a dataset of problem descriptions $x_{i}$ and the corresponding unit tests $ut_{i}$ that can be used for sampling and testing new solutions from the CLM, denoted $CLM_{base}$ henceforth. Since fine-tuning the model on the limited solutions $y_{i}$ would lead to overfitting, we leverage its extensive pretraining to generate a multitude of diverse solutions to obtain additional training data. We sample 100 solutions from $CLM_{base}$ for each problem description with multinomial sampling due to its lower computational cost. A temperature of $t=0.6$ is applied to achieve a balance between functional correctness and diversity, resulting in non-uniform runtimes.
Algorithm 1 Timing module algorithm.
$s\in solutions$
1:$CoV\leftarrow\infty$Repeat $\triangleright$ up to 1K times
[…]
4:$times.append(runtime)$EndFor
[…]
7$\triangleright$ stable runtime was not obtained
[…]
2.2 Annotation
The solutions are automatically evaluated for functional correctness and runtime. While the former can be achieved by simply executing a solution with its corresponding unit tests, the latter requires additional steps for obtaining stable runtime measurements, see Algorithm 1. Each solution $s$ is executed 50 times to determine its functional correctness (passed/failed) and $runtime$ in nanoseconds. We obtain $\mu$ and $\sigma$, then calculate the coefficient of variation $CoV$. A measurement is deemed stable and accepted if $CoV\leq 0.1$ (usually much lower). Otherwise, we repeat the loop up to 1K times. In the unlikely scenario that a stable $runtime$ could not be obtained, we set $passed=False$ (mark solution as failed). In order to further increase the reliability of $runtime$ measurements, we execute Algorithm 1 five times (in a separate process) and average the results. Lastly, we remove problems $x_{i},y_{i},ut_{i}$ which do not have at least two passing and one failed solution to ensure that optimisation can be enhanced with our Dynamic Solution Selection (2.4) during training. The statistics of the final dataset $D_{train}$ are shown in Table 3.
2.3 Optimisation
In this step, the model is fine-tuned on $D_{train}$ to bias $CLM_{base}$ towards generating more functionally correct and runtime-efficient solutions. Although several methods for preference data optimisation exist (Yuan et al., 2023; Zhao et al., 2023; Liu et al., 2024; Azar et al., 2023; Ethayarajh et al., 2024; Hong et al., 2024), we opt for DPO due to its simplicity and wide adoption. We also benchmark SFT due to its widespread use in prior work.
Supervised Fine-Tuning
We train $CLM_{base}$ on $D_{train}$ using TOP-$N$of the fastest solutions where $N\in\{25,100\}$, which means that the diversity of runtimes grows as $N$ increases. Henceforth, models optimised with the top 25% of fastest solutions are denoted as $SFT_{25}$ and CLMs trained with all (including the slowest) solutions as $SFT_{100}$.
$\mathcal{L}_{\mathrm{SFT}}\left(\pi_{\theta}\right)=-\mathbb{E}_{\left(x,y\right)\sim D}\left[\log\pi_{\theta}\left(y\mid x\right)\right]$ (1)
Direct Preference Optimisation
Aiming to avoid the complexity and instability of reinforcement learning, DPO (Rafailov et al., 2024) aligns models to preference data with a simple classification loss, shown in Equation 2.
$\mathcal{L}_{\mathrm{DPO}}\left(\pi_{\theta};\pi_{\mathrm{ref}}\right)=-\mathbb{E}_{\left(x,y_{w},y_{l}\right)\sim\mathcal{D}}\\ \log\sigma\left(\beta\log\frac{\pi_{\theta}\left(y_{w}\mid x\right)}{\pi_{\mathrm{ref}}\left(y_{w}\mid x\right)}-\beta\log\frac{\pi_{\theta}\left(y_{l}\mid x\right)}{\pi_{\mathrm{ref}}\left(y_{l}\mid x\right)}\right)$ (2)
We investigate the effectiveness of the following configurations of code preference pairs:
  • Quick versus Slow: Choose quick & slow solutions according to the annotated runtime. We denote such models as $DPO_{QvS}$.
  • Passed versus Failed: Choose passed & failed pairs according to the annotated functional correctness, denoted as $DPO_{PvF}$.
  • All: Choose all preference pairs from the Quick vs. Slow and Passed vs. Failed configurations. We denote such models as $DPO_{All}$.
Figure 2: The $pass@k$ scores for MBPP and HumanEval averaged across model sizes for a high-level overview. Models optimised via DPO consistently show higher functional correctness compared to Base and SFT for all $k$.
Figure 3: The median runtime and code length of generated solutions for MBPP and HumanEval, averaged across model sizes. Values shown are the percentage changes relative to Base, i.e. >0 is slower or longer than Base, <0 is faster or shorter. The best DPO models achieve a reduced runtime compared to SFT models as well as the very competitive Base models. A significant reduction in code length (10% - 20%) is observed across both datasets.
2.4 Dynamic Solution Selection
Training data is typically fixed at the start of training and remains static throughout Tunstall et al. (2023); Luo et al. (2023); Xu et al. (2023); Wang et al. (2023); Yuan et al. (2024). Our approach takes advantage of the multitude of code solutions from the sampling step (2.1) to dynamically select preference pairs during training. To that end, we randomly choose a new preference pair ($y_{w},y_{l}$) for each problem $x_{i}$ from $D_{train}$ at the start of the epoch for DPO configurations. For SFT, we randomly choose any working solution ($y_{w}$) at the start of each epoch for a comparable configuration. This reduces overfitting by presenting prompts with multiple completions. Note that we utilise dynamic solution selection by default in our framework.
3 Results
In this section, we define the evaluation metrics and present the results of our proposed framework at varying scales. We also provide a qualitative analysis to support our findings. Detailed implementation notes are provided in Appendix A.
Figure 4: The $pass@1$, $pass@10$ and $pass@100$ scores for MBPP and HumanEval as the number of parameters increases. A significant improvement over competitive Base and SFT models can be observed for DPO configs.
3.1 Evaluation Metrics
Functional Correctness
is evaluated by sampling 100 solutions per problem via multinomial sampling and a temperature of $t=0.6$. Following  Chen et al. (2021), we measure functional correctness using $pass@k$, where $k\in\{1,10,100\}$.
Code Efficiency
improvements can be a challenge to capture accurately. Using Algorithm 1, we measure efficiency using runtime (the median of all working solutions). Since the runtime of a failed program is undefined, we remove problems for which models have no working solutions to compare CLMs on the same subset of solved problems. Doing so ensures a fair comparison between models. Table 2 shows that this subset increases as CLMs get larger and more ‘code-competent’.
Model MBPP HumanEval
StarCoder-1B 40.60% 30.49%
StarCoder-3B 48.40% 46.95%
CodeLlama-7B 55.60% 73.71%
CodeLlama-13B 60.40% 79.27%
Table 2: Intersection of problems between $Base$, $SFT$, and $DPO$ models with at least one working solution.
Code Length
does not necessarily correlate with code efficiency as shorter outputs may abstract away the complexities of their implementations. Note that Code-Optimise does not explicitly fine-tune CLMs for code length. However, we are still interested in determining if our preference optimisation results in code that is both faster (execution savings) and shorter (inference savings). The subset of working solutions in Table 2 is again used to measure code length, which is the median number of characters of all working solutions.
3.2 Functional Correctness
Figure 2 shows the $pass@k$ scores for MBPP and HumanEval, averaged over all model sizes. The individual $pass@k$ scores are shown in Figure 4. We observe that models optimised via DPO consistently demonstrate higher functional correctness relative to the baseline (Base) and SFT on both datasets. The effect is even larger on in-domain data, particularly with lower $k$. The DPO models perform similarly on MBPP with $DPO_{PvF}$ being the best overall on HumanEval. SFT models show a marginal improvement for $k=1$ but no improvement (or a small decrease) at higher $k$. We therefore conclude that DPO is a more suitable fine-tuning paradigm for our self-generated code preference data as it is better able to leverage the learning signals (quick versus slow and passed versus failed).
Figure 5: The runtimes for MBPP and HumanEval as model size increases. Values shown are the percentage changes relative to Base, i.e. >0 means slower than Base, <0 means faster. On average, DPO models show a greater runtime reduction on in-domain rather than out-of-domain data. SFT models exhibit inconsistent scaling patterns.
Figure 6: Code lengths for MBPP and HumanEval as model sizes increase. Values shown are the percentage changes relative to Base, i.e. >0 means longer than Base, <0 means shorter. DPO models consistently produce shorter sequences across both datasets. SFT models generate significantly longer code, particularly the larger CLMs.
3.3 Code Efficiency
The runtimes and lengths of generated solutions are plotted in Figure 3 as a percentage change from the baseline (a value $<0$ means faster or shorter than the baseline while $>0$ means slower or longer). Once again, values are averaged across model sizes for a high-level overview. Individual model scores are shown in Figures 5 and 6, respectively. In preliminary analysis, we observed that baseline CLMs were already capable of generating solutions with low-complexity. However, $DPO_{QvS}$ and $DPO_{All}$ models manage to further decrease runtimes on in-domain data by up to 6% and up to 3% on the out-of-domain data. SFT models generally increase runtimes across both datasets. In terms of code length, the best DPO models reduce the median number of characters by up to 48% on MBPP and 23% on HumanEval while SFT models tend to generate significantly longer solutions. This is particularly evident with $SFT_{100}$, which uses all code solutions for training, including the slowest, which tend to be longer. SFT does not appear to be particularly suitable for optimising runtime or code length with our preference data as any baseline biases for generating longer code can be reinforced. In summary, Code-Optimise induces a reduction in runtime for faster code execution while also outputting shorter solutions, resulting in lower inference costs and improved response times.
Figure 7: The $pass@1$ scores for StarCoder-1B without (Static) and with (Dynamic) solution selection (DSS). DSS benefits every model, especially DPO configs. More $pass@k$ scores can be found in Figure 12 of the Appendix.
3.4 Model Scaling
Figures 4, 5 and 6 show the evolution of functional correctness, runtimes and lengths of generated solutions as the number of trainable parameters increases (1B - 13B). Analysing $pass@1$ in Figure 4, we can see that larger DPO models achieve a more significant improvement over the baseline and SFT, particularly for in-domain problems. Somewhat surprisingly, functional correctness for HumanEval (out-of-domain) improves at a faster rate than MBPP (up to 7B parameters). In Figure 5, we observe that as the DPO models increase in size, their runtimes relative to the baseline remain consistent. The $DPO_{PvF}$ configuration tends to average somewhat slower runtimes as this setup only optimises for correctness thus sacrificing efficiency. We can also see a consistent pattern of increased runtimes for all SFT models. On HumanEval, on the other hand, runtimes for different model sizes are much less predictable. However, on average, our best configuration $DPO_{QvS}$ does show an improvement over the already competitive baseline CLMs.
The effect on code length generalises very well to out-of-domain problems, particularly for larger CLMs, see Figure 6. In fact, we find a clear trend for all DPO models and for both datasets that shows reduced code lengths of up to 48% in-domain and up to 23% out-of-domain. The SFT models increase the lengths in all cases, especially at larger model sizes. As was the case with runtimes, this is akin to reinforcing its biases towards more verbose code as the preference data is self-generated.
3.5 Qualitative Analysis
Figure 8 shows several solutions to a typical programming problem taken from MBPP that gives a more tangible form to our results. More examples for HumanEval and MBPP can be found in the Appendix (Figures 10 and 11). Following a manual inspection of dozens of generated solutions from each configuration, the efficiency improvements generally come from two main sources: (1) brevity: the model outputs only essential code (no function calls, unit tests, comments, etc.), which saves generation time for auto-regressive LMs and (2) complexity: the code is simplified and uses faster routines, relative to the baseline, which saves resources when it is executed. The SFT models tend to sacrifice brevity the most as their complexity is similar to the baseline. Figures 8, 10, and 11 show several examples of this, e.g. adding function calls to the newly generated solution, possibly with import statements and/or expected inputs or outputs [^1] Adding test cases to code is generally considered good programming practice. However, for the purpose of pure efficiency (our case), this can potentially add significant cost.. This is in line with the observation from Figure 3 where the SFT models appear to be more verbose and biased towards longer outputs. The DPO models tend to produce solutions with a somewhat lower complexity and a better unit test coverage. Further analysis suggests that HumanEval solutions generated by baseline LMs are quite competitive and usually more runtime-efficient than MBPP baseline solutions. We posit that this may be due to the more comprehensive task descriptions in HumanEval, which include input-output pairs. Among DPO models, we do not observe a clear winner in qualitative analysis although $DPO_{QvS}$ is the best setup in terms of aggregate results.
Input prompt: def test_duplicate(arraynums):
 """Write a function to find whether a given array of integers contains any duplicate element."""
Base Solution: for i in range(0, len(arraynums)-1):
 for j in range(i+1, len(arraynums)):
 if arraynums[i] == arraynums[j]:
 return True
return False
SFT25 Solution: # Your code here
for i in range(len(arraynums)):
 for j in range(i+1, len(arraynums)):
 if arraynums[i] == arraynums[j]:
 return True
return False
array = [1, 2, 3, 1]
print(test_duplicate(array))
# Write a function to find the frequency of the elements present in a given array.
SFT100 Solution: # Code here
arraynums.sort()
for i in range(len(arraynums)-1):
 if arraynums[i] == arraynums[i+1]:
 return True
return False
if __name__ == ’__main__’:
 arraynums = [1, 2, 3, 4, 5, 1]
 print(test_duplicate(arraynums))
DPOQvS Solution: return len(arraynums) != len(set(arraynums)) DPOPvF Solution: # Write your code here
if len(arraynums) == len(set(arraynums)):
 return False
return True
DPOAll Solution: nums = set(arraynums)
return len(nums) != len(arraynums)
Figure 8: Example solutions for an MBPP problem with additional use cases found in Figures 10 and 11.
3.6 Importance of Solution Selection
Our core methodology for creating high-quality code preference data enables us to dynamically select unique pairs for each prompt at the start of a new epoch. Since we train all models for 30 epochs, CLMs can potentially be exposed to many unique combinations of code completions. Figure 7 shows $pass@1$ scores for StarCoder-1B improving with dynamic solution selection compared to static pairs randomly assigned at the beginning of training, commonly practiced in related work. The benefits are somewhat more pronounced for DPO, our preferred optimisation method, compared to SFT. Across different $k$ in $pass@k$, models consistently benefit from our dynamic solution selection, additionally seen in Figure 12 in the Appendix.
4 Related Work
Distilled Supervised Fine-Tuning
was applied to code generation models for the sole purpose of improving functional correctness. CLMs such as MagiCoder Wei et al. (2023) and WizardCoder Luo et al. (2023) query large proprietary models to provide the necessary training data Cui et al. (2023); Xu et al. (2023). However, such approaches may not be permissible in many development settings due to financial, legal and/or privacy reasons.
Reinforcement Learning
Le et al. (2022); Wang et al. (2022); Gorinski et al. (2023) can overcome the shortcomings of supervised fine-tuning by effectively propagating the negative rewards for dysfunctional code. However, RL algorithms typically come with additional complexity and instability. Recently, Rafailov et al. (2024) proposed Direct Preference Optimisation (DPO) as an alternative to Reinforcement Learning from Human Feedback for aligning language models with human preferences Tunstall et al. (2023). DPO serves as a form of offline reinforcement learning that directly optimises on a given set of trajectories without the need for a separate reward model. We should note that the aforementioned RL approaches to code synthesis only consider functional correctness and not the runtime of generated solutions.
Code Efficiency Optimisation
was previously proposed by Shypula et al. (2023) as a code editing/repair task where slow-running code had to be edited to achieve a faster runtime. Models were optimised on a newly curated dataset augmented by synthetic test cases through various methods of prompting and fine-tuning. However, the greatly reduced runtimes came at a significant cost to functional correctness. In many configurations, the model edits reduce performance by up to 30% with ‘smaller’ CLMs (7B, 13B) suffering a larger degradation. We hypothesise that this may be due to a) overfitting the single runtime objective (in contrast with our work where the aim is to optimise both correctness and runtime) and b) the removal of failed programs from the dataset leading the CLMs to struggle with semantics of correct versus incorrect code. We opt to not compare directly with this work as their method is specifically curated to the code editing task where an already functionally correct but inefficient program is assumed as input. On the contrary, we aim to produce programming solutions from scratch that are both functionally correct and runtime/inference efficient.
5 Conclusions
We have introduced Code-Optimise, a lightweight framework for improving functional correctness and runtime via self-generated code preference data optimisation (quick versus slow and passing versus failing solutions). Our experiments have shown several benefits: 1) functional correctness is significantly improved, particularly for smaller models, 2) dynamic solution selection during training provides an additional benefit by reducing overfitting, 3) runtime is reduced by up to 6% for MBPP and up to 3% for HumanEval over strong baseline CLMs, lowering the cost of code execution, 4) code length is significantly shortened, up to 48% for MBPP and up to 23% for HumanEval, which reduces inference cost and improves response times. We hope that our insights as well as our novel dataset will stimulate further exciting research in this area.
6 Limitations
Timing the execution of short programs accurately is challenging and despite our best efforts, the runtime measurements could probably be improved further with additional software engineering efforts. This would also provide a cleaner and more stable learning signal for Code-Optimise, which could potentially improve results. While our methodology is highly data-efficient, using only $\sim$200 open-source prompts for training data generation, obtaining additional high-quality problems (free from proprietary/licensing issues) may potentially yield better results. Other code-related tasks that may be amenable to optimisation for improved runtime/inference could potentially benefit from our methodology and as such may be investigated outside of the scope of this paper. While we conducted all experiments using Python, we acknowledge that less popular/similar programming languages should also be investigated in follow-up work.
Appendix A Implementation Details
A.1 Dataset
MBPP
The Mostly Basic Programming Problems introduced by Austin et al. (2021) consists of 974 crowd-sourced Python programming challenges. Each problem comprises a description, an example code solution and a few automated test cases. The dataset contains training, validation and test splits. We utilise the training and validation splits for optimisation, while the test split serves as the in-domain test data distribution.
HumanEval
Chen et al. (2021) comprises 164 Python programming challenges. The function signatures, docstrings, example solutions and several unit tests were handwritten for each problem. We leverage HumanEval as our out-of-domain test set as the descriptions in MBPP do not contain any unit tests and the writing style of HumanEval problems does not follow a consistent format. This helps us evaluate robustness to handwritten prompts.
A.2 Training
We use the StarCoder Li et al. (2023) and CodeLlama Rozière et al. (2024) families of models in our experiments. We opt for the pretrained (base) versions with sizes of 1B and 3B for StarCoder and 7B and 13B for CodeLlama, hosted on HuggingFace Wolf et al. (2020). During training, we fine-tune each model using a total of 30 epochs and select the best model based on the lowest validation loss. We use a learning rate of $5e^{-7}$ with a linear scheduler, a 10% warm-up, and a maximum sequence length of 2048 tokens.
Appendix B Supplementary Experiments
B.1 Additional Qualitative Examples
In Figures 10 and 11, we present additional qualitative examples from each configuration.
B.2 Additional Ablation Scores
In Figure 12, we present additional $pass@10$ and $pass@100$ scores for MBPP and HumanEval of StarCoder-1B by ablating the solution selection.
B.3 Fastest Solution Analysis
Shypula et al. (2023) introduce the Best@k metric, which considers only the fastest solution given $k$ samples. We show the results of our optimisation using this non-standard metric as an additional analysis. We set $k=100$ (all generated solutions), which is the basis of all our experiments. In Figure 9, we note that DPO models produce faster solutions not only on in-domain problems, but also out-of-domain, between 2% and 5% faster. $DPO_{PvF}$ once again has the higher runtime as its objective is to optimise only functional correctness. The fastest solutions from the SFT models are generally slower on both MBPP and HumanEval. Note that Best@k may overestimate the runtime improvements by only considering the fastest solution. Hence, we utilise the median of all working solutions as a less biased evaluation in our experiments.
Figure 9: The $best@100$ scores for MBPP and HumanEval, averaged across model sizes. Values shown are the percentage change to Base, i.e. >0 is slower than Base, <0 is faster. Considering the fastest solution for each problem, DPO models show the best runtimes.
Input prompt: def has_close_elements(numbers, threshold):
"""Check if in a list of numbers, are any two numbers closer to each other than threshold.
 has_close_elements([1.0, 2.0, 3.0], 0.5) == False
 has_close_elements([1.0, 2.8, 3.0, 4.0, 5.0, 2.0], 0.3) == True"""
Base Solution: for index, number in enumerate(numbers):
 for other_number in numbers[index+1:]:
 if abs(number - other_number) < threshold:
 return True
return False
if __name__ == ’__main__’:
 import doctest
 doctest.testmod()
SFT25 Solution: for i, number1 in enumerate(numbers[:-1]):
 for number2 in numbers[i+1:]:
 if abs(number1 - number2) < threshold:
 return True
return False
if __name__ == "__main__":
 import doctest
 doctest.testmod()
 # print(has_close_elements([1.0, 2.8, 3.0, 4.0, 5.0, 2.0], 0.4))
 # print(has_close_elements([1.0, 2.8, 3.0, 4.0, 5.0, 2.0, 6.0], 0.4))
<five additional test cases truncated>
SFT100 Solution: for i, number1 in enumerate(numbers[:-1]):
 for number2 in numbers[i+1:]:
 if abs(number1 - number2) < threshold:
 return True
return False
if __name__ == "__main__":
 import doctest
 doctest.testmod()
 # print(has_close_elements([1.0, 2.0, 3.0], 0.5))
 # print(has_close_elements([1.0, 2.8, 3.0, 4.0, 5.0, 2.0], 0.3))
<six additional test cases truncated>
DPOQvS Solution: return any(
 abs(numbers[i] - numbers[j]) < threshold
 for i in range(len(numbers))
 for j in range(i + 1, len(numbers))
DPOPvF Solution: for i, x in enumerate(numbers[:-1]):
 for y in numbers[i+1:]:
 if abs(x - y) < threshold:
 return True
return False
DPOAll Solution: for i in range(len(numbers) - 1):
 for j in range(i + 1, len(numbers)):
 if abs(numbers[i] - numbers[j]) < threshold:
 return True
return False
Figure 10: Example solutions for a problem in HumanEval.
Input prompt: def remove_Occ(s,ch):
 """ Write a python function to remove first and last occurrence of a given character from the string. """
Base Solution: if ch in s:
 s = s.replace(ch, ’’, 1) # remove first occurrence
if ch in s:
 s = s.replace(ch, ’’, 1) # remove second occurrence
return s
SFT25 Solution: if ch in s:
 return s.replace(ch, "", 1).replace(ch, "", 1)
else:
 return s
# s = "geeksforgeeks"
# ch = "e"
# print(remove_Occ(s, ch))
# Write a python function to remove all the vowels from the string.
SFT100 Solution: l = []
for i in s:
 if i != ch:
 l.append(i)
return "".join(l)
if __name__ == ’__main__’:
 s = input().strip()
 ch = input().strip()
 result = remove_Occ(s, ch)
 print(result)
 # Given a string, return a version without the first and last char of the string. The string may be any length -- including 0 -- so you must handle the case of an empty string.
DPOQvS Solution: while ch in s:
 s = s.replace(ch,"",1)
return s
DPOPvF Solution: return s.replace(ch,’’,2) DPOAll Solution: return s.replace(ch,"",2).replace(ch,"",-1)
Figure 11: Example solutions for a problem in MBPP.
(a) $pass@10$
(b) $pass@100$
Figure 12: The $pass@10$ and $pass@100$ scores for StarCoder-1B without (Static) and with (Dynamic) solution selection (DSS). Performance improves on both metrics and distributions with DSS.
Appendix C Complete Results
C.1 Sampling
In Table 3, we show the functional correctness of the CLMs in the sampling stage of Code-Optimise.
Model Split Pass@1 Pass@10 Pass@100
StarCoder-1B Train 14.00 34.50 55.20
Validation 12.20 31.70 48.90
StarCoder-3B Train 19.50 44.30 61.70
Validation 19.20 42.50 57.80
CodeLlama-7B Train 25.80 54.00 70.10
Validation 23.40 50.30 68.90
CodeLlama-13B Train 28.80 58.20 71.60
Validation 24.60 52.90 66.70
Table 3: Functional correctness of the CLMs during sampling.
C.2 Optimisation
In Tables 4, 5, 6, and 7, we tabulate the full model performance of the CLMs on the test sets. The CoV is shown beside each runtime.
Model Pass@1 Pass@10 Pass@100 Time Length
$Base$ 11.80 31.70 49.80 114338 $\pm$ 0.021 155
$SFT_{25}$ 17.90 34.40 47.60 104690 $\pm$ 0.012 238
$SFT_{100}$ 16.80 34.20 47.00 169536 $\pm$ 0.017 252
$DPO_{QvS}$ 17.10 36.10 52.80 109051 $\pm$ 0.018 144
$DPO_{PvF}$ 16.90 36.90 54.00 118418 $\pm$ 0.019 181
$DPO_{All}$ 16.90 36.40 53.20 103588 $\pm$ 0.021 152
(a) MBPP
Model Pass@1 Pass@10 Pass@100 Time Length
$Base$ 12.00 24.30 39.00 150930 $\pm$ 0.017 124
$SFT_{25}$ 14.20 24.30 39.00 157975 $\pm$ 0.027 180
$SFT_{100}$ 13.90 24.50 40.20 154395 $\pm$ 0.020 175
$DPO_{QvS}$ 14.20 27.30 42.10 143259 $\pm$ 0.013 125
$DPO_{PvF}$ 14.30 28.10 45.70 147980 $\pm$ 0.034 146
$DPO_{All}$ 13.70 27.10 42.10 232759 $\pm$ 0.012 132
(b) HumanEval
Table 4: Model performance on MBPP and HumanEval of StarCoder-1B.
Model Pass@1 Pass@10 Pass@100 Time Length
$Base$ 16.90 40.00 55.00 113760 $\pm$ 0.016 158
$SFT_{25}$ 23.40 41.80 55.20 115834 $\pm$ 0.011 171
$SFT_{100}$ 22.40 41.60 55.20 119675 $\pm$ 0.035 198
$DPO_{QvS}$ 23.80 46.10 59.80 112395 $\pm$ 0.008 162
$DPO_{PvF}$ 23.90 45.50 60.20 116529 $\pm$ 0.017 185
$DPO_{All}$ 23.40 45.30 60.20 103726 $\pm$ 0.012 149
(a) MBPP
Model Pass@1 Pass@10 Pass@100 Time Length
$Base$ 17.20 36.80 61.00 143806 $\pm$ 0.012 162
$SFT_{25}$ 19.20 38.80 56.10 149743 $\pm$ 0.017 172
$SFT_{100}$ 19.40 38.60 56.10 152948 $\pm$ 0.022 190
$DPO_{QvS}$ 21.00 42.90 67.70 151401 $\pm$ 0.011 170
$DPO_{PvF}$ 21.50 44.30 70.10 153620 $\pm$ 0.013 181
$DPO_{All}$ 20.50 42.30 66.50 147823 $\pm$ 0.014 161
(b) HumanEval
Table 5: Model performance on MBPP and HumanEval of StarCoder-3B.
Model Pass@1 Pass@10 Pass@100 Time Length
$Base$ 21.40 48.50 65.20 105313 $\pm$ 0.012 196
$SFT_{25}$ 25.40 48.40 62.00 124000 $\pm$ 0.058 372
$SFT_{100}$ 24.30 49.10 62.60 110982 $\pm$ 0.010 435
$DPO_{QvS}$ 28.60 52.00 66.80 108925 $\pm$ 0.013 141
$DPO_{PvF}$ 30.20 52.10 66.20 109783 $\pm$ 0.006 129
$DPO_{All}$ 29.10 52.30 66.60 108992 $\pm$ 0.016 129
(a) MBPP
Model Pass@1 Pass@10 Pass@100 Time Length
$Base$ 25.10 55.00 79.30 646547 $\pm$ 0.004 188
$SFT_{25}$ 26.80 55.00 82.90 509264 $\pm$ 0.004 256
$SFT_{100}$ 26.40 54.10 82.30 496296 $\pm$ 0.006 304
$DPO_{QvS}$ 28.20 60.30 84.80 562279 $\pm$ 0.005 159
$DPO_{PvF}$ 30.10 64.00 86.60 639553 $\pm$ 0.003 166
$DPO_{All}$ 28.70 61.20 85.40 646486 $\pm$ 0.002 160
(b) HumanEval
Table 6: Model performance on MBPP and HumanEval of CodeLlama-7B.
Model Pass@1 Pass@10 Pass@100 Time Length
$Base$ 23.70 52.50 67.60 118418 $\pm$ 0.009 223
$SFT_{25}$ 28.80 53.70 66.20 112624 $\pm$ 0.006 348
$SFT_{100}$ 26.70 52.80 66.00 126165 $\pm$ 0.004 523
$DPO_{QvS}$ 33.50 56.40 70.60 110390 $\pm$ 0.008 116
$DPO_{PvF}$ 34.10 55.50 69.00 110427 $\pm$ 0.018 126
$DPO_{All}$ 32.80 56.20 69.20 110679 $\pm$ 0.008 122
(a) MBPP
Model Pass@1 Pass@10 Pass@100 Time Length
$Base$ 27.80 62.70 87.20 497649 $\pm$ 0.015 187
$SFT_{25}$ 30.00 62.70 85.40 560336 $\pm$ 0.005 238
$SFT_{100}$ 27.90 61.00 82.90 532856 $\pm$ 0.006 375
$DPO_{QvS}$ 32.60 67.40 88.40 513372 $\pm$ 0.005 145
$DPO_{PvF}$ 33.20 68.00 88.40 528546 $\pm$ 0.008 157
$DPO_{All}$ 31.90 66.70 86.00 520788 $\pm$ 0.003 141
(b) HumanEval
Table 7: Model performance on MBPP and HumanEval of CodeLlama-13B.